the measure of a supplement of a angle is 24 degrees more than three times the measure of the original angle. find the measures of the angles
Answer by macston(5194)

.
x=original angle; s=supplement=3x+24
.
x+s=180
x+(3x+24)=180
4x=156
x=39
ANSWER 1: The original angle is 39 degrees.
.
s=3x+24
s=3(39)+24
s=117+24=141
ANSWER 2: The supplement is 141 degrees.